# Intrinsic viscous liquid dynamics

### v1

@@ Abstract
When liquids are cooled, their dynamics are slowed, and if crystallization is avoided, they will solidify into an amorphous structure referred to as a glass. Experiments show that chemically distinct glass-forming liquids have universal features of the spectrum and temperature dependence of the main structural relaxation. We introduce Randium, a generic energetically coarse-grained model of viscous liquids, and demonstrate that the intrinsic dynamics of viscous liquids emerges. These results suggest that Randium belongs to a universal class of systems whose dynamics capture the essential physics of viscous liquid relaxation, bridging microscopic molecular models and coarse-grained theoretical descriptions.
preprint: Draft version
@@ I Introduction
Molecular motion becomes slower as liquids are cooled. If crystallization is bypassed, the system solidifies into a disordered structure termed a glass [Dyre2007]. At the glass transition temperature, the viscosity of the liquid becomes so large that it ceases to flow. Various experiments have suggested that chemically distinct glass-forming liquids exhibit generic dynamics in their relaxation spectra and temperature dependence of structural relaxation. This work is directly motivated by recent depolarized dynamic light-scattering investigations by Böhmer, Pabst, Weigl, Helbling, Richter, Gabriel, Zeißler, and Blochowicz [Pabst2021, Bhomer2025], which provide striking evidence confirming a longstanding hypothesis that the spectral shape associated with structural relaxation in molecular liquids can be collapsed onto a common shape [Kohlrausch1854, Cole1941, Davidson1950, Williams1970, Jonscher1977, Jonscher1977, Derrida1980, Derrida1981, Dyre1988, Monthus1996, Bierwirth2017, Zhang2017, Niss2018, Dyre2000, Pabst2021, Denny2003, Dyre2024, Bhomer2025, Hasyim2024, Ong2024, Lam2025, Railton2025]. In this paper, we aim to provide a framework for explaining the generic relaxation of highly viscous liquids.
@@ II Randium
Dynamics is defined through Monte Carlo simulations with nearest-neighbor swap attempts, employing Boltzmann’s acceptance criterion. The physical interpretation of a neighbor particle swap is a local fundamental collective motion from one inherent state to another of the fine-grained system. This dynamics ensure that rearrangements are local, and that back jumps are likely. The unit of time is defined as one attempt per particle of the model. Conveniently, the system can be equilibrated at low temperatures with unphysical swaps of particle identities in an ensemble coupled to a bath with a distribution of polydisperse particles, resulting in $N=M$. Both types of dynamics can be implemented using a parallelizable algorithm, allowing for efficient calculations on a graphics processing unit. Below we present results with local particle swaps for a system size of $L=192$ (N=36 864N=36\,864) using between two and 512 independent initial configurations. For this system size, one million swap attempts per particle on an NVIDIA GeForce RTX 4070 take about 5 minutes.
@@
Figure 3: (a) The overlap order-parameter $Q(t/\tau)$ and (b) $\log(1-Q(t/\tau))$ as a function of scaled time. The orange dashed curve indicates a universal curve that $Q(t)$ approaches at intermediate and long times.
@@
Figure 7: (a) Reduced length scale ($\xi/\xi_{0}$) of dynamical heterogeneity at $t=\tau$ (Fig. 6) estimates as a exponential fit, $\exp(-r/\xi)$, to the ”spin-spin” correlation function, $G(r)=\langle\sigma_{i,j}\sigma_{i,j+r}\rangle$ where $\sigma_{i,j}=-1$ if the site is unchanged, and $\sigma_{i,j}=+1$ otherwise (inspired by analysis of the 2D Ising model [Shrock2022]). $\xi/\xi_{0}$ is shown agains the reduced inverse temperatur, $\beta/\beta_{g}$. (b) The reduced length scale ($\xi/\xi_{0}$) as a function of relaxation time $\log(\tau)$. The black dashed line is a $\tau\propto\exp(\xi)$ fit, suggesting that that $\xi/\xi_{0}=6$ at the glass-transition temperture ($\tau_{g}=10^{14}$).
@@ III Results
To monitor dynamics, we define the overlap, $O(t)$, as the fraction of sites that are occupied by the same particle after a time interval $t$. Let $Q(t)=\langle O(t)\rangle$ be the overlap function averaged over initial configurations. Fig. 2(a) shows $Q(t)$ for inverse temperatures ($\beta$’s) ranging from zero (dark red) to 2.0 (dark blue). At high temperatures, the relaxation is nearly exponential (black dashed), whereas at low temperatures, it follows a stretched exponential with an exponent of $\tfrac{1}{2}$ (red dashed): $A\exp(-\sqrt{t/t_{0}})$, where $A\simeq 0.98$. In the frequency domain, this corresponds to a minimum slope of the main relaxation peak of $\tfrac{1}{2}$, consistent with dielectric experiments [Nielsen2009]. At the lowest temperatures, a plateau develops from particle backjumps, as shown in Fig. 2(b) by plotting $\log(Q(t)-1)$. This explains why $A$ is slightly less than one in an empirical fit to a stretched exponential.
@@ IV Discussion
How does Randium compare to other proposed explanations of generic viscous liquid dynamics? Historically, the first descriptions are empirical approaches such as fits to a stretched exponential [Kohlrausch1854, Williams1970] in the time-domain, or the Cole-Cole fit in the frequency-domain [Cole1941, Davidson1950].
More theoretically founded approaches include kinetic facilitation models [Lam2025], spin-glass models [Edwards1975, Sherrington1975, Derrida1980, Derrida1981, Hopfield1982, Parisi1979, Parisi1980, Parisi1983, Sherrington2025, Dahlberg2025], energetic barrier and trap models [Dyre1988, Bouchaud1992, Monthus1996, Diezemann1997, Monthus2003, Schrder2008], elastic models [Dyre2006, Vasin2022, Hasyim2024], and the recently proposed Hyper-sphere model [Railton2025]. Like many of these approaches, Randium builds on the idea of an intrinsic energy landscape put forward by Goldstein in 1969 [Goldstein1969]. In particular, the distinguishable-particle lattice (DPL) model by Lam and coworkers [Zhang2017, Ong2024, Lee2025], and the lattice-gas on a random energy landscape in three dimensions [Pedersen2006] are closely related. Like Randium, these models are defined as particles on a lattice – unlike Randium, dynamics are defined as particles moving into a void, like in the kids’ toy 15-Puzzle [Ratner1990]. The motivation for this dynamics is string-like motions seen in computational studies of atomic glass-formers [Donati1998, Schroder2000]. In Randium, the swap of two particles is interpreted as a fundamental transition between local minima, which, for the atomic model, are string-like motions of tens of particles [Vogel2004]. Thus, Randium can be viewed as more coarse-grained than the models of Refs. [Pedersen2006, Zhang2017, Ong2024, Lee2025] allowing for a large dynamical range. Dynamical mobile regions of the DPL model are dictated by the locations of voids, in contrast, dynamical heterogeneity is an emergent phenomenon.
@@ Acknowledgements.
The authors thank Thomas Blochowicz, Mark Ediger, Tina Hecksher, Andreas Heuer, Camille Scalliet, Walter Kob, Thomas B. Schrøder, and Jeppe C. Dyre for helpful discussions, and in particular Florian Pabst for curating and discussing experimental data.

### v2

@@ Abstract
When liquids are cooled, their dynamics are slowed, and if crystallization is avoided, they will solidify into an amorphous structure referred to as a glass. Experiments show that chemically distinct glass-forming liquids have universal features of the spectrum and temperature dependence of the main structural relaxation. We introduce Randium, a generic energetically coarse-grained model of viscous liquids, and demonstrate that the intrinsic dynamics of viscous liquids emerges. The results suggest that Randium belongs to a class of systems whose dynamics capture the essential physics of viscous liquid relaxation, bridging microscopic molecular models and coarse-grained theoretical descriptions.
@@ I Introduction
Molecular motion becomes slower as liquids are cooled. If crystallization is bypassed, the system solidifies into a disordered structure termed a glass [Dyre2007]. At the glass transition temperature, the viscosity of the liquid becomes so large that it ceases to flow. Various experiments have suggested that chemically distinct glass-forming liquids exhibit generic dynamics in their relaxation spectra and temperature dependence of structural relaxation. This work is directly motivated by recent depolarized dynamic light-scattering investigations by Pabst, Böhmer, Weigl, Helbling, Richter, Zourchang, Walther, Gabriel, Zeißler, and Blochowicz [Pabst2021, Bhomer2025], which provide striking evidence confirming a longstanding hypothesis that the spectral shape associated with structural relaxation in molecular liquids can be collapsed onto a common shape [Kohlrausch1854, Cole1941, Davidson1950, Williams1970, Jonscher1977, Jonscher1977, Derrida1980, Derrida1981, Dyre1988, Monthus1996, Bierwirth2017, Zhang2017, Niss2018, Dyre2000, Pabst2021, Denny2003, Dyre2024, Bhomer2025, Hasyim2024, Ong2024, Lam2025, Railton2025]. In this paper, we aim to provide a framework for explaining the generic relaxation of highly viscous liquids.
@@ II Randium
Dynamics is defined through Monte Carlo simulations with nearest-neighbor swap attempts, employing Boltzmann’s acceptance criterion. The physical interpretation of a neighbor particle swap is a local fundamental collective motion from one inherent state to another of the fine-grained system. This dynamics ensure that rearrangements are local, and that back jumps are likely. The unit of time is defined as one attempt per particle of the model. Conveniently, the system can be equilibrated at low temperatures with unphysical swaps of particle identities in an ensemble coupled to a bath with a distribution of polydisperse particles, resulting in $N=M$. Both types of dynamics can be implemented using a parallelizable algorithm, allowing for efficient calculations on a graphics processing unit. Below we present results with local particle swaps for a system size of $L=192$ (N=36 864N=36\,864) using between two and 512 independent initial configurations. For this system size, one million swap attempts per particle on an NVIDIA GeForce RTX 4070 take about 5 minutes. An Python implementation is available at the DOI 10.5281/zenodo.17554510.
@@
Figure 3: (a) The overlap order-parameter $Q(t/\tau)$ and (b) $\log(1-Q(t/\tau))$ as a function of scaled time. The orange dashed curve indicates a universal curve that $Q(t)$ approaches at intermediate and long times. For comparison, the red dashed is a stretch exponential: $Q=0.98\exp(\sqrt{t/t_{0}})$.
@@
Figure 7: (a) Reduced length scale ($\xi/\xi_{0}$; $\xi_{0}=\xi(\beta=0)$) of dynamical heterogeneity at $t=\tau$ (Fig. 6) estimates as a exponential fit, $\exp(-r/\xi)$, to the ”spin-spin” correlation function, $G(r)=\langle\sigma_{i,j}\sigma_{i,j+r}\rangle$ where $\sigma_{i,j}=-1$ if the site is unchanged, and $\sigma_{i,j}=+1$ otherwise (inspired by analysis of the 2D Ising model [Shrock2022]). $\xi/\xi_{0}$ is shown agains the reduced inverse temperatur, $\beta/\beta_{g}$. (b) The reduced length scale ($\xi/\xi_{0}$) as a function of relaxation time $\log(\tau)$. The black dashed line is a $\tau\propto\exp(\xi)$ fit, suggesting that that $\xi/\xi_{0}=6$ at the glass-transition temperture ($\tau_{g}=10^{14}$).
@@ III Results
To monitor dynamics, we define the overlap, $O(t)$, as the fraction of sites that are occupied by the same particle after a time interval $t$. Let $Q(t)=\langle O(t)\rangle$ be the overlap function averaged over initial configurations. Fig. 2(a) shows $Q(t)$ for inverse temperatures ($\beta$’s) ranging from zero (dark red) to 2.0 (dark blue). At high temperatures, the relaxation is nearly exponential (black dashed), whereas at low temperatures, it resembles a stretched exponential with an exponent of $\tfrac{1}{2}$ (red dashed): $A\exp(-\sqrt{t/t_{0}})$, where $A\simeq 0.98$. In the frequency domain, this corresponds to a minimum slope of the main relaxation peak of $\tfrac{1}{2}$, consistent with dielectric experiments [Nielsen2009]. At the lowest temperatures, a plateau develops from particle backjumps, as shown in Fig. 2(b) by plotting $\log(Q(t)-1)$. This explains why $A$ is slightly less than one in an empirical fit to a stretched exponential.
@@ IV Discussion
How does Randium compare to other proposed explanations of generic viscous liquid dynamics? Historically, the first descriptions are empirical approaches such as fits to a stretched exponential [Kohlrausch1854, Williams1970] in the time-domain, or the Cole-Cole fit in the frequency-domain [Cole1941, Davidson1950]. More theoretically founded approaches include kinetic facilitation models [Lam2025], spin-glass models [Edwards1975, Sherrington1975, Derrida1980, Derrida1981, Hopfield1982, Parisi1979, Parisi1980, Parisi1983, Sherrington2025, Dahlberg2025], energetic barrier and trap models [Dyre1988, Bouchaud1992, Monthus1996, Diezemann1997, Monthus2003, Schrder2008], elastic models [Dyre2006, Vasin2022, Hasyim2024], and the recently proposed Hyper-sphere model [Railton2025]. Like many of these approaches, Randium builds on the idea of an intrinsic energy landscape put forward by Goldstein in 1969 [Goldstein1969]. In particular, the distinguishable-particle lattice (DPL) model by Lam and coworkers [Zhang2017, Ong2024, Lee2025], and the lattice-gas on a random energy landscape in three dimensions [Pedersen2006] are closely related. Like Randium, these models are defined as particles on a lattice – unlike Randium, dynamics are defined as particles moving into a void, like in the kids’ toy 15-Puzzle [Ratner1990]. The motivation for this dynamics is string-like motions seen in computational studies of atomic glass-formers [Donati1998, Schroder2000]. Randium has no adjustable parameter associated with void concentration. In Randium, the swap of two particles is interpreted as a fundamental transition between local minima, which, for the atomic model, are string-like motions of tens of particles [Vogel2004]. Thus, Randium can be viewed as more coarse-grained than the models of Refs. [Pedersen2006, Zhang2017, Ong2024, Lee2025] allowing for a large dynamical range. Dynamical mobile regions of the DPL model are dictated by the locations of voids, in contrast, dynamical heterogeneity is an emergent phenomenon in Randium.
@@ V Concluding remarks
Data and code availability
Data and an Python implementation of Randium are available from Zenodo at DOI 10.5281/zenodo.17554510 [data]. The repository contain additional figures showing the determination of fit-parameters in Fig. 5, and fits to determine the $\xi$ length-scales presented in Fig. 7.
@@ Acknowledgements.
The author thank Thomas Blochowicz, Mark Ediger, Tina Hecksher, Andreas Heuer, Camille Scalliet, Walter Kob, Nicholas Bailey, Thomas B. Schrøder, and Jeppe C. Dyre for helpful discussions, and in particular Florian Pabst for curating and discussing experimental data.